Solve the system by graphing both equations on the same axes. Give the ordered pair for the point of intersection.
x+y=6
3x-y=-2

Answer by mananth(16946)

x+y=6
3x-2y=-2
Find the x & y intercept points by plugging x=0,y=0 in both equations.
equation I
when x= 0 y= 6 ,( 0 , 6 )
when y = 0 x= 6 ,( 6 , 0 )
Equation 2
when x= 0 y= -1 ( 0 , -1 )
when y = 0 x= -0.67 ( -0.67 , 0 )


(1,5)